OpenStudy (anonymous):

3/48 is what part of 1/12

OpenStudy (asnaseer):

suppose the question was: 3 is what part of 12? how would you work out the answer?

OpenStudy (anonymous):

3 is 4th part of of 12 now?

OpenStudy (asnaseer):

correct - so you divided 3 by 12 to get:\[3\div12=\frac{1}{4}\]to say that 3 is a quarter of 12. Use the same principal for your question.

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

not compltly can you explain please

OpenStudy (asnaseer):

ok, lets look at another example - 5 is what part of 10? to find out, we just divided 5 by 10 to get:\[5\div10=\frac{5}{10}=\frac{1}{2}\]so we can say 5 is half of 10.

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

can you give me example of any fraction question/

OpenStudy (asnaseer):

ok, lets do: 1/4 is what part of 1/2? again to find out we just divide as follows:\[\frac{1}{4}\div\frac{1}{2}=\frac{1}{4}\times\frac{2}{1}=\frac{1}{2}\]so we can say 1/4 is a half of 1/2

OpenStudy (anonymous):

it means 3/48 ÷ 1/12 = 3/48 × 12/1 = 1

OpenStudy (anonymous):

am i right?

OpenStudy (asnaseer):

how did you get 3/48 x 12/1 = 1?

OpenStudy (anonymous):

as you gave example

OpenStudy (anonymous):

please solve my question

OpenStudy (asnaseer):

the example I gave was: